OpenStudy (anonymous):

Tony has 6 dollars in change in his car. Every time he drives to visit his grandma, he pays $1.25 in tolls. Which graph represents the change Tony has in his car?

OpenStudy (anonymous):

these are the graphs that go with it

OpenStudy (anonymous):

help asap please

OpenStudy (texaschic101):

graph 4 is your answer take 2 points on the line...and if your slope is -1.25, then it is the right graph....the slope is negative because you are paying money (4,1) and (0,6) slope = (6 - 1) / (0 - 4) slope = 5/-4 slope = -5/4 which equals - 1.25

OpenStudy (anonymous):

thank you i understand now :)